Find the intercept of the graph of this equation. 8x-3y=24
Answer by unlockmath(1688)

Hello,
If you want the y intercept then let's get the equation in the form of
y=mx+b (Hint: b is the y intercept)
8x-3y=24 Subtract 8x and divide by -3 to get:
y=-8/3(x)-3
OK, now we know that -3 on the y axis is where the line passes through.
Make sense?
RJ
www.math-unlock.com